Anil and Babs are playing a game in which they take turns to throw a fair six-sided die. The winner is the first player to throw the same number as the other player has thrown immediately before. Anil throws first.

  1. (a)

    Find exactly the probability that Anil wins the game.

  2. (b)

    Given that Babs wins the game, find exactly the probability that she wins on her second throw.

Worked solution

(a)

1. 分析每轮掷骰子获胜与继续的概率

游戏从第 22 次掷骰子开始,每次掷骰子的玩家若掷出与前一人相同的点数即获胜,其概率为 16\frac{1}{6};若不同则游戏继续,概率为 1−16=561 - \frac{1}{6} = \frac{5}{6}。
第 11 次掷骰子由 Anil 进行,此时无人掷过,无法获胜。

2. 列出 Anil 获胜的概率表达式并求和

Anil 只能在第 3,5,7,…,2k+13, 5, 7, \dots, 2k+1 次掷骰子时获胜:

  • 第 33 次获胜(Anil 的第 22 次投掷):56×16=536\frac{5}{6} \times \frac{1}{6} = \frac{5}{36}
  • 第 55 次获胜(Anil 的第 33 次投掷):(56)3×16\left(\frac{5}{6}\right)^3 \times \frac{1}{6}
  • 第 2k+12k+1 次获胜:(56)2k−1×16\left(\frac{5}{6}\right)^{2k-1} \times \frac{1}{6}

Anil 获胜的总概率为首项 a=536a = \frac{5}{36},公比 r=(56)2=2536r = \left(\frac{5}{6}\right)^2 = \frac{25}{36} 的无穷等比级数之和:
P(Anil 获胜)=a1−r=5361−2536=5361136=511P(\text{Anil 获胜}) = \frac{a}{1 - r} = \frac{\frac{5}{36}}{1 - \frac{25}{36}} = \frac{\frac{5}{36}}{\frac{11}{36}} = \frac{5}{11}

Answer

511\frac{5}{11}

(b)

1. 求出 Babs 获胜的总概率

因为博弈以概率 11 结束,且获胜者非 Anil 即 Babs,所以:
P(Babs 获胜)=1−P(Anil 获胜)=1−511=611P(\text{Babs 获胜}) = 1 - P(\text{Anil 获胜}) = 1 - \frac{5}{11} = \frac{6}{11}

2. 计算 Babs 在其第 2 次投掷获胜的概率

Babs 的第 22 次投掷是整场游戏的第 44 次投掷。前 33 次均未结束游戏,且第 44 次与第 33 次点数相同:
P(Babs 在第 2 次投掷获胜)=(56)3×16=1251296P(\text{Babs 在第 2 次投掷获胜}) = \left(\frac{5}{6}\right)^3 \times \frac{1}{6} = \frac{125}{1296}

3. 利用条件概率公式计算最终概率

由条件概率公式:
P(Babs 在第 2 次投掷获胜∣Babs 获胜)=P(Babs 在第 2 次投掷获胜)P(Babs 获胜)=1251296611=1251296×116=2751296P(\text{Babs 在第 2 次投掷获胜} \mid \text{Babs 获胜}) = \frac{P(\text{Babs 在第 2 次投掷获胜})}{P(\text{Babs 获胜})} = \frac{\frac{125}{1296}}{\frac{6}{11}} = \frac{125}{1296} \times \frac{11}{6} = \frac{275}{1296}

Answer

2751296\frac{275}{1296}

AI analysis — may contain mistakes.

Worked this one out on paper? Snap your steps and Gigle's AI math checker will circle the step where it went wrong.

More questions on this topic

  1. 1A bag contains one green disc, rr red discs and 2r2r blue discs, where r>1r>1. Tim and Sim play a game by taking turns t…
  2. 2On the first day of last month, the temperature of a machine in a manufacturing plant was found to be 90∘C90^\circ\text{C}…
  3. 3An arithmetic progression has first term aa and common difference dd, where aa and dd are non-zero. The first, third…
  4. 4A finite sequence {an}\{a_n\} has 50 terms and is such that an+1=an+0.15a_{n+1} = a_n + 0.15 for n=1,2,3,…,49n = 1,2,3,\ldots,49. Another infi…
  5. 5In a computer simulation, each player uses an identical pail to pour water repeatedly into his assigned tank, which is i…